Poplar products is evaluating a piece of machinery. The cost of the machinery is $1,700,000. The machinery will be depreciated over five years. At the end of the project, the equipment will be worth 25% of its original value. The tax rate is 35%. What is the after-tax proceeds at the end of the project?
Answer #1

Calculation of after-tax proceeds at the end of the project

Cost of Machinery =17,00,000

So salvage value = 17,00,000*25%=425,000

So after tax proceeds = 425,000*(1-.35) =276,250